# ARTalk: Speech-Driven 3D Head Animation via Autoregressive Model

### v2

@@ Abstract
Speech-driven 3D facial animation aims to generate realistic lip movements and facial expressions for 3D head models from arbitrary audio clips. Although existing diffusion-based methods are capable of producing natural motions, their slow generation speed limits their application potential. In this paper, we introduce a novel autoregressive model that achieves real-time generation of highly synchronized lip movements and realistic head poses and eye blinks by learning a mapping from speech to a multi-scale motion codebook. Furthermore, our model can adapt to unseen speaking styles using sample motion sequences, enabling the creation of 3D talking avatars with unique personal styles beyond the identities seen during training. Extensive evaluations and user studies demonstrate that our method outperforms existing approaches in lip synchronization accuracy and perceived quality. Demos and codes are available at https://xg-chu.site/project_artalk/.
1 Introduction
@@
Figure 1: We present ARTalk, a framework for speech-driven 3D facial motion generation. Our method learns a mapping from speech to a multi-scale motion codebook, enabling the real-time generation of realistic and diverse animation sequences.
Speech-driven 3D facial animation has attracted significant attention in both academia and industry due to its broad applications in virtual reality, game animation, film production, and human-computer interaction. This task focuses on generating natural and synchronized facial expressions, particularly lip movements, along with realistic head motions directly from speech input. In recent years, many exploratory methods have significantly enhanced the performance of speech-driven motion generation tasks. However, generating natural and human-like motions remains challenging due to the inherently complex many-to-many mapping between speech and facial or head movements. Capturing these intricate relationships is crucial for creating realistic animations.
Some existing methods, such as (Richard et al., 2021; Xing et al., 2023; Sung-Bin et al., 2024; Nocentini et al., 2024), employ autoregressive models to predict deformation offsets on meshes for speech-driven motion generation. However, due to the high complexity of meshes in 3D space, these approaches often fail to fully capture the many-to-many relationships between speech and motion. This limitation frequently results in overly smooth non-lip facial expressions that lack realism. Recent works leveraging diffusion models (Ho et al., 2020; Tevet et al., 2023; Stan et al., 2023; Sun et al., 2024) have demonstrated impressive results in generative tasks, including speech-driven motion. For instance, FaceDiffuser (Stan et al., 2023) employs diffusion to generate motions directly on meshes, while DiffPoseTalk (Sun et al., 2024) generates facial expressions and head motions using diffusion on FLAME (Li et al., 2017) blendshapes, achieving natural results. However, these methods involve computationally expensive diffusion processes and cannot perform real-time and low-latency motion generation. Meanwhile, multi-scale autoregressive models (Tian et al., 2024) have shown success in image autoregressive generation tasks, offering fast and high-quality results. However, these models are designed for single multi-scale samples instead of temporal motions, and thus cannot be directly applied to speech-to-motion generation tasks.
@@
To address the challenges of achieving high-quality, natural, and fast speech-driven motion generation while overcoming these limitations, we propose a novel method based on sliding temporal windows and multi-scale autoregression. Our method divides speech into time windows for generation, which is important for low-latency action generation. We extend the VQ autoencoder to encode multi-scale motion codes across two consecutive time windows, enabling effective handling of temporal sequence data. Additionally, we design a novel autoregressive generator that conditionally generates multi-scale motion codes based on the speech input of the current time window and previously generated motions. This design allows our autoregressive model to perform high-quality multi-scale generation within each time window while maintaining temporal continuity across windows. Furthermore, we integrated a transformer-based style encoder to extract style codes from sample motion segments, enabling our model to generate stylized motions that reflect unique individual characteristics while reducing the complexity of the many-to-many mapping between speech and motion. These innovations enable our method to achieve personalized and high-fidelity lip motions, facial expressions and head motion generation, making it suitable for a wide range of speech-driven downstream tasks.
@@
2 Related Work
Table 1: Comparison across methods, including style adaptation during inference (w/ Style), head pose generation (w/ Pose), and real-time capability (Real-time). ARTalk (ours) is the only method to achieve all three features, demonstrating its comprehensive advantages over baseline methods.
@@ 2 Related Work
Figure 2: ARTalk involves two separated parts. (a) shows our temporal multi-scale VQ autoencoder. It encodes motion sequences into multi-scale token maps $[M_{k_{1}},M_{k_{2}},...,M_{K}]$ using a shared codebook and causal masking on temporal. (b) shows The ARTalk Causal Transformer, where training uses ground truth tokens with a block-wise causal attention mask, and inference autoregressively predicts motion tokens conditioned on speech features and last scale tokens and last time window motions.
2.1 Speech-Driven 3D Facial Animation
Research on audio-driven 3D motion generation has been an area of interest for decades, with methods evolving significantly over time. Early approaches (Taylor et al., 2012; Xu et al., 2013; Edwards et al., 2016) relied primarily on procedural methods, which segmented speech into phonemes and mapped them to predefined visemes through handcrafted rules. Although procedural methods provide explicit control over the generated animations, they typically require complex parameter tuning and fail to capture the diversity and complexity of real-world speaking styles. As a result, procedural methods struggled to deliver animations that appeared both natural and adaptable to varying speech dynamics. In recent years, learning-based approaches (Fan et al., 2022; Xing et al., 2023; Sun et al., 2024; Daněček et al., 2023; Peng et al., 2023b; Yang et al., 2024) have advanced rapidly, addressing these limitations and enabling more natural and expressive facial animations. Additionally, some approaches (Yi et al., 2022; Ye et al., 2024; Zhang et al., 2023; Tan et al., 2024) focus on directly generating talking head videos without explicitly modeling motion. However, this limits their ability to integrate with motion-driven downstream applications, restricting their broader applicability.
2.2 Autoregressive 3D Facial Motion Generation
@@
2.3 Diffusion based 3D Facial Motion Generation
@@ 3.1 Preliminaries
3.2 Temporal Multi-scale VQ Autoencoder
@@
$\mathbf{h}^{(l)}$ $=\mathrm{Interp}\bigl{(}\mathrm{Quant}(\mathbf{r}^{(l-1)}),\,k_{l}\bigr{)},$ (1)
@@
$L_{\text{recon}}=\|\hat{M}-M\|_{1}+w_{\text{lips}}\|\hat{V}_{\text{lips}}-V_{\text{lips}}\|^{2}+\|\hat{V}-V\|^{2}$ (2)
@@
$L_{\text{vel}}$ $=\|(\hat{V}_{1:}-\hat{V}_{:-1})-(V_{1:}-V_{:-1})\|^{2},$ (3)
$L_{\text{smooth}}$ $=\|\hat{V}_{2:}-2\hat{V}_{1:-1}+\hat{V}_{:-2}\|^{2}.$ (4)
@@
$L_{\text{VQ}}=L_{\text{recon}}+\lambda_{\text{vel}}L_{\text{vel}}+\lambda_{\text{smooth}}L_{\text{smooth}}+L_{\text{cb}},$ (5)
@@
3.3 Speech-to-Motion Autoregressive Model
@@
$p(\{\mathbf{z}_{T}^{(l)}\}_{l=1}^{L}\mid\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s})\;=\;\prod_{l=1}^{L}p\!\bigl{(}\mathbf{z}_{T}^{(l)}\mid\mathbf{z}_{T}^{(<l)},\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s}\bigr{)},$ (6)
@@
$L_{\mathrm{AR}}$ $=\;-\sum_{T}\sum_{l=1}^{L}\log\,p(\mathbf{z}_{T}^{(l)\!\text{(gt)}}\mid\mathbf{z}_{T}^{(<l)\!\text{(gt)}},\mathbf{z}_{T-1}^{(L)\!\text{(gt)}},\mathbf{a}_{T},\mathbf{s}).$ (7)
@@
Figure 3: Comparison of runtime efficiency and performance across different methods.
4 Experiments
@@
Table 2: Quantitative results on the TFHP (Sun et al., 2024) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on TFHP, while all other methods were trained or fine-tuned on TFHP.
@@
Table 3: Quantitative results on the VOCASET-Test (Cudeiro et al., 2019) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on VOCASET. It is worth noting that our method is not trained or fine-tuned on VOCASET.
@@
4.1 Quantitative Results
Based on previous studies (Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Nocentini et al., 2024; Sun et al., 2024), we adopted two quantitative metrics, the lip vertex error (LVE) (Richard et al., 2021) and the upper face dynamic deviation (FDD) (Xing et al., 2023), to evaluate the generated facial motions. LVE calculates the maximum L2 error of all lip vertices for each frame, evaluating the largest deviation between predicted and ground truth lip positions. FDD calculates the standard deviation of the motion of each upper facial vertex over time between predictions and ground truth, evaluating the consistency of upper facial motion, which is closely related to speaking styles. Additionally, we use a similar metric as  Sun et al. (2024), mouth opening distance (MOD), to more accurately assess the stylistic similarity of mouth opening movements. MOD measures the average difference in mouth opening region between predictions and ground truth. Compared to LVE, MMD focuses more on the similarity of mouth opening styles and is less sensitive to temporal lip synchronization. For the partitioning of the lip area and upper face, we used the mask partitioning officially provided by FLAME (Li et al., 2017). The lip region contains 254 points, while the upper face, including the eye region and forehead, contains 884 points.
@@ 4.1 Quantitative Results
Figure 4: Qualitative comparison with existing methods (all head poses fixed). The first four rows are from the TFHP dataset, and the last two rows are from the VOCASET dataset. Our method shows better alignment with the ground truth in expression style, mouth dynamics, and lip synchronization. Additional results are available in the supplementary materials and demo videos.
4.2 Qualitative Results
In Figure 4, we present a qualitative comparison between our method and other baseline approaches. Our method demonstrates excellent lip synchronization, accurately capturing various phonetic elements. Furthermore, the generated results exhibit realistic facial expressions and mouth opening, closely matching the style of the ground truth. Notably, our method also generates realistic blinking and head movements, which are implicitly learned and encoded within the motion codebook. Additional qualitative evaluation results are available in the supplementary videos.
Table 4: User study results.
@@ 4.4 Ablation study
Figure 5: Qualitative results of head pose. When certain words are stressed or when accents occur, the model produces nodding motions similar to human behavior. Additionally, during transitions between different speech styles or sentence breaks, the head naturally turns, reflecting realistic conversational dynamics.
@@
Table 5: Ablation results on TFHP dataset.
@@ 5 Discussion and Conclusion
Acknowledgements
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. In addition, this work was also partially supported by JST SPRING, Grant Number JPMJSP2108.
Appendix A Reproducibility
A.1 More Data Processing Details of VOCASET
For evaluation, we also re-tracked the FLAME (Li et al., 2017) parameters of the VOCASET (Cudeiro et al., 2019) dataset. This is necessary because the original data of VOCASET registers head motion directly onto FLAME meshes rather than FLAME parameters, and also introduces details such as hair that cannot be represented by the FLAME parameters. Since our method and DiffPoseTalk (Sun et al., 2024) works in the FLAME parameter space, we reprocessed VOCASET accordingly. Specifically, we obtained the original VOCASET videos, resampled them to 25 fps, and used EMICA (Feng et al., 2021; Zielonka et al., 2022; Danecek et al., 2022) to extract a 100-dimensional shape vector, a 50-dimensional expression vector, and a 6-dimensional pose vector for evaluation.
For mesh-based baseline methods, we used meshes reconstructed from FLAME parameters instead of the original VOCASET meshes. While this introduces minor differences—such as the absence of hair and subtle expression details—these variations have minimal impact on our primary evaluation, which focuses on lip movements.
A.2 More Implementation Details
Our model consists of two main components: a multi-scale encoder-decoder and an autoregressive module.
The multi-scale encoder-decoder adopts a transformer-based architecture. Encoder and decoder both consist of 8 layers, each with 8 attention heads and a hidden dimension of 512. The encoder outputs a 64-dimensional codebook representation, while the decoder reconstructs the original expression and pose parameters. During multi-scale quantization, an additional fully connected layer is used in the upsampling process to recover lost information, maintaining an input and output dimension of 64. The codebook consists of 256 vectors, each with a dimensionality of 64.
The autoregressive module is a transformer-based autoregressive model. It has an input and hidden dimension of 768, a conditional vector dimension of 768, and a depth of 12 layers with 12 attention heads. At the input stage, both the codebook features and the audio encoding features are projected to a 768-dimensional space. The audio conditions are incorporated into the transformer layers using adaptive instance normalization (AdaIN) (Huang & Belongie, 2017). After generation through the transformer, a fully connected layer maps the output to a 256-dimensional space, followed by a softmax function to compute the probability distribution over the codebook entries.
We use the AdamW (Loshchilov, 2017) optimizer with an initial learning rate of $1\times 10^{-4}$, trained for 50,000 iterations. A linear decay learning rate scheduler is applied, gradually reducing the learning rate to $1\times 10^{-5}$ over the course of training.
A.3 More Evaluation Details
We generally follow the approach of CodeTalker (Xing et al., 2023) to compute LVE and FFD, and additionally compute the mean mouth opening distance (MMD). Since existing works rarely specify the exact selection of mouth and upper face vertices, we adopt the official FLAME region masks.  [^1] FLAME_masks.pkl downloaded from http://s2017.siggraph.org. Specifically, the lip/mouth region consists of 254 vertices from the “lips” area, while the upper face region includes 884 vertices from the “eye_region” and “forehead” areas.
Appendix B Preliminaries of 3DMM
We leverage the widely used 3D morphable model (3DMM), FLAME (Li et al., 2017), known for its geometric accuracy and versatility. Due to its realistic rendering capabilities and flexibility, FLAME has been widely adopted in applications such as facial animation, avatar creation, and facial recognition. We use FLAME as our representation for facial motion, construct a multi-scale codebook using FLAME parameters and learn speech-driven autoregression within this codebook, effectively leveraging the priors embedded in FLAME. This approach provides two key advantages: (1) it reduces the high-dimensional complexity of modeling a large number of mesh vertices, and (2) it enables seamless integration with downstream tasks that utilize FLAME-based representations (Chu et al., 2024; Deng et al., 2024; Chu & Harada, 2024; Xu et al., 2024).
The FLAME model represents the head shape as follows:
$TP(\hat{\beta},\hat{\theta},\hat{\psi})=\bar{T}+BS(\hat{\beta};S)+BP(\hat{\theta};P)+BE(\hat{\psi};E),$ (8)
where $\bar{T}$ represents the template head avatar mesh, $BS(\hat{\beta};S)$ is the shape blend-shape function to account for identity-related shape variation, $BP(\hat{\theta};P)$ models jaw and neck pose to correct deformations that cannot be fully explained by linear blend skinning, and the expression blend-shapes $BE(\hat{\psi};E)$ capture facial expressions such as eye closure and smiling.
Appendix C More Discussions on ARTalk
C.1 The Choice of Audio Encoder
The performance of the audio encoder has a significant impact on the overall model. In this work, we use the HuBERT (Hsu et al., 2021) encoder for experiments. However, HuBERT is primarily designed for longer audio comprehension, which may not be ideal for our windowed input setting. For instance, when the input is as short as 8 frames (320 ms), HuBERT struggles to effectively capture phonetic features.
To address this, we experiment with Mimi (Défossez et al., 2024), an audio encoder better suited for streaming input and short audio segments. As shown in Table 6, Mimi outperforms HuBERT when the input length is short, whereas HuBERT performs better on longer audio segments. We attribute this to Mimi’s design, which prioritizes capturing fine-grained phonetic details in short audio sequences. However, its lower feature dimensionality limits its representational capacity for longer segments compared to HuBERT. Ultimately, the selection of the audio encoder should be guided by the specific requirements of the application, balancing the need for short-term phonetic precision and long-term representational capacity.
Table 6: Different audio encoder performance on TFHP dataset.
[…]
Motion Clip Length 8 + MIMI 11.01 22.79 2.14
Motion Clip Length 8 + HuBERT 11.73 24.89 2.25
Motion Clip Length 25 + MIMI 10.33 19.65 1.99
Motion Clip Length 25 + HuBERT 10.20 19.00 1.97
Motion Clip Length 50 + MIMI 10.10 19.16 1.95
Motion Clip Length 50 + HuBERT 9.78 18.03 1.89
Motion Clip Length 100 + MIMI 9.97 18.48 1.93
Motion Clip Length 100 + HuBERT 9.34 18.15 1.81
C.2 Performance of Multi-scale VQ Autoencoder
The overall performance of the model is inherently influenced by the VQ autoencoder, which serves as the upper bound for the final model’s performance. Here in Table 7, we present the performance of VQ autoencoder trained with different clip lengths. We did not observe a significant impact of clip length on the performance of the VQ autoencoder. For the autoregressive generation results under different segment length settings, please refer to the Table 5 in the main paper.
Table 7: VQ Autoencoder performance on TFHP dataset.
[…]
Motion Clip Length 8 1.96 8.81 0.43
Motion Clip Length 25 1.97 9.57 0.44
Motion Clip Length 50 2.19 9.54 0.47
Motion Clip Length 100 2.08 8.97 0.45
Figure 6: The interface of our user study. Users evaluate each video based on three criteria: lip synchronization, expression naturalness, and style consistency. The first two are judged by comparing method A and B, while style consistency is assessed with reference to the ground truth. One of the videos (A or B) is generated by our method, and the other by a baseline method, with their order randomized.
Appendix D User Study Details
We collected a total of 28 survey responses, with each participant answering 84 questions (corresponding to 84 pairwise comparison trials), for a total of 2352 comparisons. Among these, 63 comparisons were conducted on the TFHP dataset and 21 on the VOCASET dataset. To mitigate potential biases caused by random or preferential selections, we randomized the display order of our method and the baseline in each trial. Which means in each comparison, one of the videos (A or B) was generated by our method, but the assignment was randomized.
For each comparison, users were asked three questions. The first two questions follow prior studies (Fan et al., 2022; Xing et al., 2023), assessing the quality of lip synchronization and the perceived naturalness of expressions. Additionally, we introduce a third question to evaluate the consistency of the generated animation with the ground truth style. The questions presented to users were as follows: Which lip sync performs better? Which looks more natural in expression? Which looks more consistent in style with the ground truth? All three questions are single-choice, requiring users to select either A or B.
Figure 7: We reconstruct the avatar using GAGAvatar (Chu & Harada, 2024) and drive it with our ARTalk model, enabling speech-driven dynamic avatar generation.
Figure 8: The interface of our user study.
Appendix E Integration with Downstream Tasks
Our method can be integrated into various FLAME-based downstream applications, enabling a wider range of use cases. While this extends beyond the primary scope of this paper, we also demonstrate its application in GAGAvatar (Chu & Harada, 2024). As shown in Figure 7, our method can generate control signals for GAGAvatar, effectively transforming it into a speech-driven one-shot dynamic avatar reconstruction method.
Appendix F More Qualitative Results
We present additional qualitative comparison results in Figure 8. However, we strongly recommend referring to the supplementary video for a more comprehensive comparison of different methods.

### v3

@@ Abstract.
Speech-driven 3D facial animation aims to generate realistic lip movements and facial expressions for 3D head models from arbitrary audio clips. Although existing diffusion-based methods are capable of producing natural motions, their slow generation speed limits their application potential. In this paper, we introduce a novel autoregressive model that achieves real-time generation of highly synchronized lip movements and realistic head poses and eye blinks by learning a mapping from speech to a multi-scale motion codebook. Furthermore, our model can adapt to unseen speaking styles, enabling the creation of 3D talking avatars with unique personal styles beyond the identities seen during training. Extensive evaluations and user studies demonstrate that our method outperforms existing approaches in lip synchronization accuracy and perceived quality. Demos and codes are available at https://xg-chu.site/project_artalk/.
Speech-driven motion generation, Head animation, Autoregressive models
journal: TOGccs: Computing methodologies Computer graphics
1. Introduction
@@
Figure 1. We present ARTalk, a framework for speech-driven 3D facial motion generation. Our method learns a mapping from speech to a multi-scale motion code, enabling the real-time generation of realistic and diverse animation sequences.
[…]
A concept pipeline of our method.
The rapid evolution of large language models (Achiam et al., 2023; Team et al., 2023; Guo et al., 2025) in recent years has driven significant progress in artificial intelligence, particularly in generating human-like text and facilitating complex conversational interactions. However, many such systems primarily rely on text or synthesized speech for communication, often lacking the visual embodiment necessary for achieving rich and intuitive human-computer interaction. To bridge this gap and create more engaging interactive experiences, research into digital humans has garnered widespread attention from both academia and industry. Due to the intrinsic and natural correlation between the speech and facial expressions, speech-driven 3D facial animation has emerged as a key component for shaping lifelike virtual characters. Therefore, advancements in this field are essential for promoting the application of digital humans in various scenarios such as education and entertainment.
Speech-driven 3D head motion generation aims to synthesize natural and temporally synchronized facial expressions, particularly lip movements, along with realistic head motions directly from speech input. Recent years have witnessed significant advancements in speech-driven motion generation. Autoregressive models, for instance, have been explored to predict deformation offsets on 3D meshes (Richard et al., 2021; Xing et al., 2023; Sung-Bin et al., 2024; Nocentini et al., 2024). However, the inherent complexity of representing facial motion through mesh deformations in 3D space often struggles to fully capture the intricate and non-linear mappings between speech and the resulting nuanced facial expressions. This limitation frequently leads to over-smoothed non-lip regions and a lack of fine-grained realism. More recently, diffusion models (Ho et al., 2020; Tevet et al., 2023; Stan et al., 2023; Sun et al., 2024) have demonstrated compelling generative capabilities, including in the domain of speech-driven motion. For example, FaceDiffuser (Stan et al., 2023) leverages diffusion processes to generate motion directly on 3D meshes, while DiffPoseTalk (Sun et al., 2024) utilizes diffusion on the parametric FLAME model (Li et al., 2017) blendshapes to produce natural facial expressions and head movements. Despite their impressive results in generating high-fidelity motions, these diffusion-based approaches typically suffer from high computational costs, hindering their applicability in real-time and low-latency scenarios. In parallel, multi-scale autoregressive models (Tian et al., 2024) have achieved notable success in image generation tasks, offering a balance between generation speed and output quality. However, these architectures are primarily designed for processing single, multi-scale static samples rather than sequential temporal data, making their direct application to the temporal dynamics inherent in speech-to-motion generation challenging.
@@
To overcome the limitations of existing methods and achieve high-quality, natural, and fast speech-driven motion generation, we propose a novel approach based on temporal windows and multi-scale autoregression. Our method segments the speech input into consecutive temporal windows for motion generation, a window-wise processing strategy crucial for achieving low-latency motion synthesis. Within each temporal window, we further employ a multi-scale encoding and generation strategy to capture facial expressions and head movements with rich details. For instance, for a 4-second temporal window containing 100 motion frames, we encode it into a set of multi-scale representations with varying temporal resolutions, such as encodings with lengths of [1, 5, 25, 50, 100], and adopt a similar multi-scale decoding process during generation. This design allows the model to simultaneously capture global temporal context and fine-grained details of each motion frame, resulting in more expressive motions. To minimize encoding errors and ensure temporal coherence of the generated motions, we extend a VQ codec to encode multi-scale motion codes across two consecutive temporal windows, thereby effectively modeling and transferring temporal dependencies between windows. Additionally, we design a novel autoregressive generator that can autoregressively predict the multi-scale motion codes for the current temporal window conditioned on the speech features of the current window and the previously generated motion sequence. This conditional generation mechanism enables our model to achieve high-quality multi-scale motion generation within each temporal window while maintaining smooth transitions and temporal consistency across windows. Furthermore, we integrate a Transformer-based style encoder to extract style representations from example motion clips. This allows our model to generate stylized motions with unique individual characteristics and helps to disentangle the complex many-to-many mapping between speech and motion. Through these innovations, our method can generate personalized and high-fidelity lip movements, facial expressions, and head motions, demonstrating significant potential for various speech-driven downstream applications.
@@
2. Related Work
Table 1. Comparison across methods, including style adaptation during inference (w/ Style), head pose generation (w/ Pose), and real-time capability (Real-time). ARTalk (ours) is the only method to achieve all three features, demonstrating its comprehensive advantages over baseline methods.
@@ 2. Related Work
Figure 2. ARTalk involves two separated parts. (a) shows our temporal multi-scale VQ autoencoder. It encodes motion sequences into multi-scale token maps $[M_{k_{1}},M_{k_{2}},...,M_{K}]$ using a shared codebook and causal masking on temporal. (b) shows The ARTalk Causal Transformer, where training uses ground truth tokens with a block-wise causal attention mask, and inference autoregressively predicts motion tokens conditioned on speech features and last scale tokens and last time window motions.
[…]
Method pipeline of our method.
2.1. Speech-Driven 3D Facial Animation
Research on audio-driven 3D motion generation has been an area of interest for decades, with methods evolving significantly over time. Early approaches (Taylor et al., 2012; Xu et al., 2013; Edwards et al., 2016) relied primarily on procedural methods, which segmented speech into phonemes and mapped them to predefined visemes through handcrafted rules. Although procedural methods provide explicit control over the generated animations, they typically require complex parameter tuning and fail to capture the diversity and complexity of real-world speaking styles. As a result, procedural methods struggled to deliver animations that appeared both natural and adaptable to varying speech dynamics. In recent years, learning-based approaches (Fan et al., 2022; Xing et al., 2023; Sun et al., 2024; Daněček et al., 2023; Peng et al., 2023b; Yang et al., 2024; Fan et al., 2024; Chae-Yeon et al., 2025) have advanced rapidly, addressing these limitations and enabling more natural and expressive facial animations. Additionally, some approaches (Yi et al., 2022; Ye et al., 2024; Zhang et al., 2023; Tan et al., 2024) focus on directly generating talking head videos without explicitly modeling motion. However, this limits their ability to integrate with motion-driven downstream applications, restricting their broader applicability.
2.2. Autoregressive 3D Facial Motion Generation
@@
2.3. Diffusion based 3D Facial Motion Generation
@@ 3.1. Preliminaries
3.2. Temporal Multi-scale VQ Autoencoder
@@
(1) $\mathbf{h}^{(l)}$ $=\mathrm{Interp}\bigl{(}\mathrm{Quant}(\mathbf{r}^{(l-1)}),\,k_{l}\bigr{)},$
@@
(2) $L_{\text{recon}}=\|\hat{M}-M\|_{1}+w_{\text{lips}}\|\hat{V}_{\text{lips}}-V_{\text{lips}}\|^{2}+\|\hat{V}-V\|^{2}$
@@
(3) $L_{\text{vel}}$ $=\|(\hat{V}_{1:}-\hat{V}_{:-1})-(V_{1:}-V_{:-1})\|^{2},$
(4) $L_{\text{smooth}}$ $=\|\hat{V}_{2:}-2\hat{V}_{1:-1}+\hat{V}_{:-2}\|^{2}.$
@@
(5) $L_{\text{VQ}}=L_{\text{recon}}+\lambda_{\text{vel}}L_{\text{vel}}+\lambda_{\text{smooth}}L_{\text{smooth}}+L_{\text{cb}},$
@@
3.3. Speech-to-Motion Autoregressive Model
@@
(6) $p(\{\mathbf{z}_{T}^{(l)}\}_{l=1}^{L}\mid\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s})\;=\;\prod_{l=1}^{L}p\!\bigl{(}\mathbf{z}_{T}^{(l)}\mid\mathbf{z}_{T}^{(<l)},\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s}\bigr{)},$
@@
(7) $L_{\mathrm{AR}}$ $=\;-\sum_{T}\sum_{l=1}^{L}\log\,p(\mathbf{z}_{T}^{(l)\!\text{(gt)}}\mid\mathbf{z}_{T}^{(<l)\!\text{(gt)}},\mathbf{z}_{T-1}^{(L)\!\text{(gt)}},\mathbf{a}_{T},\mathbf{s}).$
@@
Figure 3. Comparison of runtime efficiency and performance across different methods.
4. Experiments
@@
Table 2. Quantitative results on the TFHP (Sun et al., 2024) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on TFHP, while all other methods were trained or fine-tuned on TFHP.
@@
Table 3. Quantitative results on the VOCASET-Test (Cudeiro et al., 2019) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on VOCASET. It is worth noting that our method is not trained or fine-tuned on VOCASET.
@@
4.1. Quantitative Results
Based on previous studies (Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Nocentini et al., 2024; Sun et al., 2024), we adopted two quantitative metrics, the lip vertex error (LVE) (Richard et al., 2021) and the upper face dynamic deviation (FDD) (Xing et al., 2023), to evaluate the generated facial motions. LVE calculates the maximum L2 error of all lip vertices for each frame, evaluating the largest deviation between predicted and ground truth lip positions. FDD calculates the standard deviation of the motion of each upper facial vertex over time between predictions and ground truth, evaluating the consistency of upper facial motion, which is closely related to speaking styles. Additionally, we use a similar metric as  Sun et al. (2024), mouth opening distance (MOD), to more accurately assess the stylistic similarity of mouth opening movements. MOD measures the average difference in mouth opening region between predictions and ground truth. Compared to LVE, MMD focuses more on the similarity of mouth opening styles and is less sensitive to temporal lip synchronization. For the partitioning of the lip area and upper face, we used the mask partitioning officially provided by FLAME (Li et al., 2017). The lip region contains 254 points, while the upper face, including eye region and forehead, contains 884 points.
@@ 4.1. Quantitative Results
4.2. Qualitative Results
In Figure 5, we present a qualitative comparison between our method and other baseline approaches. Our method demonstrates excellent lip synchronization, accurately capturing various phonetic elements. Furthermore, the generated results exhibit realistic facial expressions and mouth opening, closely matching the style of the ground truth. Notably, our method also generates realistic blinking and head movements, which are implicitly learned and encoded within the motion codebook. In Figure 4, we also show the ability of our method on head movements. It can capture accents well and generate reasonable head movements. Additional qualitative evaluation results are available in the supplementary videos.
Table 4. User study results.
@@ 4.4. Ablation study
Figure 4. Qualitative results of head pose. When certain words are stressed or when accents occur, the model produces nodding motions similar to human behavior. Additionally, during transitions between different speech styles or sentence breaks, the head naturally turns, reflecting realistic conversational dynamics.
[…]
Qualitative comparison.
@@
Integration with Downstream Tasks Our method can be integrated into various FLAME-based downstream applications, enabling a wider range of use cases. While this extends beyond the primary scope of this paper, we also demonstrate its application in GAGAvatar (Chu and Harada, 2024). As shown in Figure 6, our method can generate control signals for GAGAvatar, effectively transforming it into a speech-driven one-shot dynamic avatar reconstruction method.
Table 5. Ablation results on TFHP dataset.